# Q*: Improving Multi-step Reasoning for LLMs with Deliberative Planning

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive capability in many nature language tasks. However, the auto-regressive generation process makes LLMs prone to produce errors, hallucinations and inconsistent statements when performing multi-step reasoning. In this paper, we aim to alleviate the pathology by introducing Q*, a general, versatile and agile framework for guiding LLMs decoding process with deliberative planning. By learning a plug-and-play Q-value model as heuristic function, our Q* can effectively guide LLMs to select the most promising next step without fine-tuning LLMs for each task, which avoids the significant computational overhead and potential risk of performance degeneration on other tasks. Extensive experiments on GSM8K, MATH and MBPP confirm the superiority of our method.
@@ 1 Introduction
Large Language Models (LLMs) have exhibited impressive capabilities in solving various reasoning tasks encoded in nature languages, including math word problems [1, 2, 3, 4, 5, 6], code generation [7, 8, 9] and planning [10, 11, 12]. Unfortunately, even the most advanced LLMs still face significant challenges and are prone to introduce errors, hallucinations and inconsistent statements as the number of reasoning steps grows due to their auto-regressive nature [13, 14]. In fact, the auto-regressive generation process of LLMs can be characterized by “System 1" [15], a mode of thought which is fast, instinctive but less accurate. Most of recent works focus on improving LLMs’ System 1 capability by (1) constructing sophisticated prompts with extensive expertise to trigger the potential capacities of LLMs without modifying their parameters [16, 17, 18, 19], (2) fine-tuning LLMs with massive task-specific corpus at the price of significant computational burdens and the potential risk of performance degeneration on other tasks [5, 6, 20, 21], or (3) training reward models to rank the candidate responses [22, 23, 4, 24].
@@
On the other hand, solving complex reasoning problems requires more in-depth, deliberative and logical thinking steps, i.e., the “System 2" mode [15]. Taking solving math word problems as an example, any incorrect intermediate reasoning step (e.g., calculation errors, mis-interpretations) can potentially lead to incorrect final answers. Prior attempts [25, 26, 27, 28] for enhancing System 2 reasoning capability includes performing deliberation with basic tree search algorithms (e.g., BFS or DFS), Monte Carlo Tree Search (MCTS) [29], and A* [30]. Nonetheless, the utility functions used in these methods often require laborious expertise to design for each specific task, which are difficult to be extended to new scenarios. Furthermore, deliberation with MCTS would require significant number of rollouts before finding high-quality responses when solving the problems with many reasoning steps, which substantially slows down the overall decoding process.
@@
In light of this, we propose Q*, a general, versatile and agile framework for improving the multi-step reasoning capability of LLMs with deliberative planning. Different from the existing deliberation methods, our method does not rely on domain knowledge to design the heuristic function. Besides, by leveraging plug-and-play Q-value models as heuristic function, our Q* can effectively solve various tasks via guiding LLMs to select the most promising next step without fine-tuning LLMs beforehand, which avoids the significant computational overhead and potential risk of performance degeneration in other tasks. Finally, Q* considers only a single step when performing deliberation, which is much cheaper than complete rollouts in MCTS. Specifically, the main contributions of our work are summarized as follows:
@@
  • We formalize the multi-step reasoning of LLMs as a Markov Decision Process (MDP) where the state is the input prompt and the reasoning steps generated so far, the action is the next step of reasoning and and the reward measures how well the task is solved.
@@
  • We present several general approaches to estimate the optimal Q-value of state-action pairs, i.e., offline reinforcement learning, best sequence from rollout and completion with stronger LLMs. It is noteworthy that our methods only need the ground truth of training problems and can be easily applied to various reasoning tasks without modification.
@@
  • We conduct extensive experiments on math reasoning and code generation tasks, demonstrating that Q* can significantly improve the multi-step reasoning capability of LLMs.
@@ 2 Related Works
Alignment is an important technique to prevent the output of LLMs deviates from human’s expectation. Supervised Fine-Tuning (SFT) is probably the most fundamental alignment approach that directly minimizes the cross-entropy loss between the output and ground-truth. Reinforcement learning from Human Feedback (RLHF) [31], on other hand, firstly learns a reward model (RM) from human preferences and then optimizes the SFT model with reinforcement learning algorithms to maximize the cumulative rewards from RM. Direct Preference Optimization (DPO) [32] aligns LLMs directly according to the ranking information from human feedback without explicitly learning RM. Recently, Aligner [33] came out as a model-agnostic alignment method by learning to re-write LLMs’ output. Compared to these methods, our Q* achieves alignment with distinct merits. Different from SFT and Aligner, Q* does not rely on massive human generated corpus which are expensive to collect; different from RLHF and DPO, our Q* does not modify the parameters of LLM, which avoid potential performance loss on other tasks.
@@
Tree-of-thoughts (ToT) [25] improves the LLMs’ reasoning capability by exploring the intermediate steps towards problem solving with basic tree-search algorithms. In the same vein, A* search and MCTS are applied to serve as a planning technique to enhance the performance of LLMs when solving challenging reasoning problems [26, 27, 28, 34]. Unfortunately, the utility function used in these methods is either constructed from LLMs’ feedback (e.g., [25, 27]), which could be highly-inaccurate in complex problems, or specific to each individual task (e.g., [28, 34]). Moreover, planning with MCTS often requires to perform costly rollout, which can significantly slow down the overall decoding process. In contrast, our Q* relies on ground-truth solely to train value model and can be easily applied to various reasoning tasks without modification. Besides, we consider only a single step each time in Q*, which is much cheaper than complete rollout in MCTS-based methods.
@@
Math reasoning and code generation require LLMs to perform multi-step reasoning on relations, quantities and logics which are inherently challenging. Current techniques include 1) prompt engineering which triggers the potential capacities of LLMs with sophisticated prompts [16, 17, 18, 19, 35, 36]. However, constructing such prompt needs extensive expertise and case-by-case tuning, which is difficult to generalize to different tasks; 2) Fine-tuning LLMs with massive math/code corpus [5, 6, 20, 21, 8, 9, 37], which usually at the price of significant computational burden and may compromise the performance on other tasks; 3) training RMs/verifiers to rank the candidate solutions without providing any guidance in intermediate steps [22, 23, 4, 24]. Differently, our Q* leverages a plug-and-play Q-value model to direct the deliberation process of LLMs, which effectively provides guidance for each intermediate step without modifying the parameters of LLMs. Moreover, by casting multi-step reasoning of LLMs as a heuristic search problem, our method can be generalized to various reasoning tasks without laborious prompt engineering.
@@ 3.1 Formulate the Multi-step Reasoning of LLMs as an MDP
Taking the question $q$ as input, the answer generation process of LLMs can be broken down into multiple reasoning steps, where the final answer sequence $\mathbf{a}$ can be treated as the concatenation of these $T$ single-step reasoning steps, formulated as $\mathbf{a}=[a_{1};a_{2};\dots;a_{T}]$. Each step can be a single line or fixed number of tokens outputted by LLMs. Under this perspective, we can conceptualize the multi-step reasoning process of LLMs as a Markov Decision Process (MDP) $\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma\rangle$, where the state $s_{t}\in\mathcal{S}$ denotes the partial reasoning trace already generated by timestep $t-1$ plus the input question (i.e., $s_{t}=[q;a_{1};\dots;a_{t-1}]$) with the special case $s_{1}=q$, the action $a_{t}\in\mathcal{A}$ denotes the next reasoning step generated by LLMs taking the current state $s_{t}$ as input, the deterministic state transition $\mathcal{T}$ from the current state $s_{t}$ to the next state $s_{t+1}$ is accomplished through a simple operation of concatenation, $\mathcal{R}$ is the reward function to measure how well the question is solved and $\gamma$ is the discount factor. The reward function is often outcome-based, i.e., it gives reward by comparing the final results with ground-truth. Finally, the policy $\pi_{\theta}$ is embodied by an LLM, which produces reasoning sequence conditioned on the input question:
@@ 3.2 A* Search
A* [30] is an important heuristic search algorithm in deliberative planning [38], multi-agent pathfinding [39], and constraint reasoning [40]. Originally, A* was proposed for finding the shortest path from source $s$ to goal $g$ in path planning problems. It associates each frontier vertex $n$ with a value $f(n)=g(n)+h(n)$, where $g(n)$ is the accumulated cost from source $s$ and $h(n)$ is a heuristic value that estimates the cost of the shortest path from $n$ to goal $g$. The algorithm adopts a best-first search strategy, i.e., in each iteration it always picks the vertex with minimum $f$-value to explore until reaching the goal. When the heuristic $h(\cdot)$ is admissible, A* guarantees to find the optimal path.
Figure 1: Overview of Q*. (a): the deliberation process of Q*. Each state is associated with a $f$-value which is the weighted sum of the aggregated utility (cf. Eq. (4)) and the heuristic value (cf. Eq. (5)). (b-d): estimating optimal Q-value with fitted-Q-iteration, rollout and completion with stronger LLMs.
@@ 4 Q*: A General, Versatile and Agile Deliberation Framework for LLMs
Most of modern LLMs generate natural languages in an auto-regressive way, i.e., predict the next token in a sequence given the previously generated tokens (cf. Eq. (1)). Therefore, when applied to multi-step reasoning problem, LLMs can potentially introduce errors, hallucinations and inconsistent statements in subsequent reasoning trace if any previous step is incorrect, which may fail to solve the problem. Indeed, given the fact that LLMs produce each token with fixed computation, there is no way to investigate more computational power to solve difficult problems. Simply put, LLMs cannot perform in-depth deliberation which is essential for solving complex reasoning problem.
We address the issue by presenting Q*, a general, versatile and agile deliberation framework based on A* to effectively guide LLMs to select the most promising next step when perform multi-step reasoning without costly fine-tuning LLMs for each task beforehand. In more detail, we cast finding the most proper reasoning sequence for a given problem as a heuristic search process, where each state $s_{t}$ is associated with a $f$-value estimating how much utility will be attained if we expand $s_{t}$:
@@
where $g(s_{t})$ is the aggregated utility from the initial state $s_{1}$, $h(s_{t})$ is the heuristic value measuring the quality of the final result derived from $s_{t}$ and $\lambda$ is a coefficient to balance $g(s_{t})$ and $h(s_{t})$.
@@
Since maximizing over all possible next reasoning steps is intractable, in practice we restrict the alternatives to the top-K steps returned by LLM, i.e., $f(s_{t})=g(s_{t})+\lambda\max_{a_{t}\in\text{top-K}(\pi_{\theta}(\cdot|s_{t}))}Q^{*}(s_{t},a_{t})$.
@@ 4.1 Estimation of Optimal Q-value
where $s_{t}=[q_{i};a_{1};\dots;a_{t-1}]$ is the partial reasoning trace by timestep $t-1$ in $\mathbf{a}_{i_{j}}$ and $\hat{y}(s_{t},a_{t})$ is the label that approximates the true optimal Q-value $Q^{*}(s_{t},a_{t})$.
@@
Staring from state-action pair $(s_{t},a_{t})$, we first perform random rollout or MCTS [27, 26] with policy $\pi_{\theta}$ for enough iterations, forming a trajectory pool $P$. After that, we use the best reasoning sequence with the highest accumulated reward to construct Q-value label:
@@ 4.2 Deliberative Planning with A*
Once obtaining proxy Q-value model $\hat{Q}$, we can plug it to Eq. (5) to compute the $f$-value of each state and perform best-first search with A*. Alg. 1 illustrates the deliberative planning process.
@@
2:$pq\leftarrow\{q\}$, $visited\leftarrow\emptyset$
3:while $pq\neq\emptyset$ do
4:     $s\leftarrow\arg\max_{s^{\prime}\in pq}f(s^{\prime})$
5:     $pq\leftarrow pq\backslash\{s\}$, $visited\leftarrow visited\cup\{s\}$
6:     if $s$ is a terminal state then return extract-answer($s$)
@@
9:         if $s^{\prime}\notin visited$ then $pq\leftarrow pq\cup\{s^{\prime}\}$
Specifically, we maintain an open list $pq$ which initially contains the input question, and a closed list $visited$ to record the visited states. Each step we pick the state $s$ with the maximum $f$-value from the open list and expand it by querying the top-K alternatives with the LLM policy $\pi_{\theta}$. After that, the open list is updated and the process repeats until the terminal state is reached. Finally, we extract the answer part of the terminal state as the result.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive capability in many natural language tasks. However, the auto-regressive generation process makes LLMs prone to produce errors, hallucinations and inconsistent statements when performing multi-step reasoning. In this paper, by casting multi-step reasoning of LLMs as a heuristic search problem, we aim to alleviate the pathology by introducing Q*, a general, versatile and agile framework for guiding LLMs decoding process with deliberative planning. By learning a plug-and-play Q-value model as heuristic function for estimating expected future rewards, our Q* can effectively guide LLMs to select the most promising next reasoning step without fine-tuning LLMs for the current task, which avoids the significant computational overhead and potential risk of performance degeneration on other tasks. Extensive experiments on GSM8K, MATH and MBPP demonstrate the superiority of our method, contributing to improving the reasoning performance of existing open-source LLMs.
@@ 1 Introduction
Large Language Models (LLMs) have exhibited impressive capabilities in solving various reasoning tasks encoded in natural languages, including math word problems [1, 2, 3, 4, 5, 6], code generation [7, 8, 9] and planning [10, 11, 12]. Unfortunately, even the most advanced LLMs still face significant challenges and are prone to introduce errors, hallucinations and inconsistent statements as the number of reasoning steps grows due to their auto-regressive nature [13, 14]. In fact, the auto-regressive generation process of LLMs can be characterized by “System 1" [15], a mode of thought which is fast, instinctive but less accurate. Most of recent works focus on improving LLMs’ “System 1” capability by (1) constructing sophisticated prompts with extensive expertise to trigger the potential capacities of LLMs without modifying their parameters [16, 17, 18, 19], (2) fine-tuning LLMs with massive task-specific corpus at the price of significant computational burdens and the potential risk of performance degeneration on other tasks [5, 6, 20, 21], or (3) training reward models to rank the candidate responses [22, 23, 4, 24].
@@
On the other hand, solving complex reasoning problems requires more in-depth, deliberative and logical thinking steps, i.e., the “System 2" mode [15]. Taking solving math word problems as an example, any incorrect intermediate reasoning step (e.g., calculation errors, mis-interpretations) can potentially lead to incorrect final answers. Prior attempts [25, 26, 27, 28] for enhancing “System 2” reasoning capability includes performing deliberation with basic tree search algorithms (e.g., BFS or DFS), Monte Carlo Tree Search (MCTS) [29], and A* [30]. Nonetheless, the utility functions used in these methods often require laborious expertise to design for each specific task, which are difficult to be extended to new scenarios. Furthermore, deliberation with MCTS would require significant number of rollouts before finding high-quality responses when solving the problems with many reasoning steps, which substantially slows down the overall decoding process.
@@
In light of this, we propose Q*, a general, versatile and agile framework for improving the multi-step reasoning capability of LLMs with deliberative planning. Different from the existing deliberation methods, our method does not rely on domain knowledge to design the heuristic function. Besides, by leveraging plug-and-play Q-value models as heuristic function, our Q* can effectively solve various tasks via guiding LLMs to select the most promising next step without fine-tuning LLMs beforehand, which avoids the significant computational overhead and potential risk of performance degeneration in other tasks. Finally, Q* considers only one single step when performing deliberation, which is much cheaper than completing rollouts in MCTS. Specifically, the main contributions of our work are summarized as follows:
@@
  • We formalize the multi-step reasoning of LLMs as a Markov Decision Process (MDP) where the state is the concatenation of input prompt and the reasoning steps generated so far, the action is the next reasoning step and the reward measures how well the task is solved.
@@
  • We present several general approaches to estimate the optimal Q-value of state-action pairs, i.e., offline reinforcement learning, the best sequence from rollouts, and completion with stronger LLMs. It is noteworthy that our methods only need the ground-truth of training problems and can be easily applied to various reasoning tasks without modification.
@@
  • We conduct extensive experiments on math reasoning and code generation tasks, demonstrating that Q* can significantly improve the multi-step reasoning capability of existing open-sourced LLMs.
@@ 2 Related Works
Alignment has become an important technique to prevent the output of LLMs deviates from human’s expectation. Supervised Fine-Tuning (SFT) is probably the most fundamental alignment approach that directly minimizes the cross-entropy loss between the output and ground-truth. Reinforcement learning from Human Feedback (RLHF) [31], on the other hand, firstly learns a reward model (RM) from human preferences and then optimizes the SFT model with reinforcement learning algorithms to maximize the cumulative rewards from RM. Direct Preference Optimization (DPO) [32] aligns LLMs directly according to the ranking information from human feedback without explicitly learning RM. Recently, Aligner [33] came out as a model-agnostic alignment method by learning to re-write LLMs’ output. Compared to these methods, our Q* achieves the goal of alignment with distinct merits. Different from SFT and Aligner, Q* does not rely on massive human annotated preference pairs which are expensive to collect; different from RLHF and DPO, our Q* does not modify the parameters of LLMs, which avoids the potential risk of performance degeneration on other tasks.
@@
Tree-of-thoughts (ToT) [25] improves the LLMs’ reasoning capability by exploring the intermediate steps towards problem solving with basic tree-search algorithms. In the same vein, A* search and MCTS are applied to serve as a planning technique to enhance the performance of LLMs when solving challenging complex reasoning problems [26, 27, 28, 34]. Unfortunately, the utility function used in these methods is either constructed from LLMs’ feedback (e.g., [25, 27]), which could be highly-inaccurate in complex problems, or specific to each individual task (e.g., [28, 34]). Moreover, planning with MCTS often requires to perform costly rollout, which can significantly slow down the overall decoding process. In contrast, our Q* sorely relies on training a Q-value model to guide LLMs to select the most promising next reasoning step and the pipeline can be easily applied to various reasoning tasks without modification. Besides, we consider only a single step each time in Q*, which is much cheaper than complete rollout in MCTS-based methods.
@@
Math reasoning and code generation require LLMs to perform multi-step reasoning on relations, quantities and logics which are inherently challenging. Current techniques include: 1) prompt engineering which triggers the potential capacities of LLMs with sophisticated prompts [16, 17, 18, 19, 35, 36]. However, constructing such prompt needs extensive expertise and case-by-case tuning, which is difficult to generalize to different tasks; 2) Fine-tuning LLMs with massive math/code corpus [5, 6, 20, 21, 8, 9, 37], which usually comes at the price of significant computational burden and may compromise the performance on other tasks; 3) training RMs/verifiers to rank the candidate solutions without providing any guidance in intermediate steps [22, 23, 4, 24]. Differently, our Q* leverages a plug-and-play Q-value model to direct the deliberation process of LLMs, which effectively provides guidance for each intermediate step without modifying the parameters of LLMs. Moreover, by casting multi-step reasoning of LLMs as a heuristic search problem, our method can be generalized to various reasoning tasks without laborious prompt engineering.
@@ 3.1 Formulate the Multi-step Reasoning of LLMs as an MDP
Taking the question $q$ as input, the answer generation process of LLMs can be broken down into multiple reasoning steps, where the final answer sequence $\mathbf{a}$ can be treated as the concatenation of these $T$ single-step reasoning steps, formulated as $\mathbf{a}=[a_{1};a_{2};\dots;a_{T}]$. Each step can be a single line or fixed number of tokens outputted by LLMs. Under this perspective, we can conceptualize the multi-step reasoning process of LLMs as a Markov Decision Process (MDP) $\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma\rangle$, where the state $s_{t}\in\mathcal{S}$ denotes the concatenation of the input question and the partial reasoning trace already generated by timestep $t-1$ (i.e., $s_{t}=[q;a_{1};\dots;a_{t-1}]$) with the special case $s_{1}=q$, the action $a_{t}\in\mathcal{A}$ denotes the next reasoning step generated by LLMs taking the current state $s_{t}$ as input, the deterministic state transition $\mathcal{T}$ from the current state $s_{t}$ to the next state $s_{t+1}$ is accomplished through a simple operation of concatenation, $\mathcal{R}$ is the reward function to measure how well the question is solved and $\gamma$ is the discount factor. The reward function is often outcome-based. That is, it gives reward by comparing the final results with ground-truth:
$\mathcal{R}(s_{t},a_{t})=\begin{cases}1&t=T\land[s_{t};a_{t}]\;\text{matches the ground-truth}\\ 0&\text{otherwise}\end{cases},$ (1)
In particular, we assign a reward of 1 if the generated code passes all test cases (for code generation tasks) or the final answer matches the ground-truth (for math reasoning tasks) which is a common practise in previous studies [4, 22]. Finally, the policy $\pi_{\theta}$ is embodied by an LLM, which produces reasoning sequence conditioned on the input question:
@@ 3.2 A* Search
A* [30] is an important heuristic search algorithm in deliberative planning [38], multi-agent pathfinding [39], and constraint reasoning [40]. Originally, A* is proposed for finding the shortest path from source $s$ to goal $g$ in path planning problems. It associates each frontier vertex $n$ with a value $f(n)=g(n)+h(n)$, where $g(n)$ is the accumulated path cost from source $s$ and $h(n)$ is a heuristic value that estimates the cost of the shortest path from $n$ to goal $g$. The algorithm adopts a best-first search strategy, i.e., in each iteration it always picks the vertex with minimum $f$-value to explore until reaching the goal. When the heuristic $h(\cdot)$ is admissible [41], A* guarantees to find the optimal path.
Figure 1: Overview of Q*. (a): the deliberation process of Q*. Each state is associated with an $f$-value which is the weighted sum of the aggregated utility (cf. Eq. (5)) and the heuristic value (cf. Eq. (6)). (b-d): estimating optimal Q-value with fitted-Q-iteration, rollout and completion with stronger LLMs.
@@ 4 Q*: A General, Versatile and Agile Deliberation Framework for LLMs
Most of modern LLMs generate natural languages in an auto-regressive way, i.e., predict the next token in a sequence given the previously generated tokens (cf. Eq. (2)). Therefore, when applied to multi-step reasoning problem, LLMs can potentially introduce errors, hallucinations and inconsistent statements in the subsequent reasoning trace if any previous step is incorrect, which may fail to solve the current problem. Indeed, given the fact that LLMs produce each token with limited computation resources, there is no way to devote more computational efforts to solve difficult problems. In short, LLMs cannot perform in-depth deliberation which is essential for solving complex multi-step reasoning problems.
We address this issue by presenting Q*, a general, versatile and agile deliberation framework based on A* to effectively guide LLMs to select the most promising next step when performing multi-step reasoning without costly fine-tuning LLMs for each task beforehand. In more detail, we cast finding the most proper reasoning sequence for a given problem as a heuristic search process, where each state $s_{t}$ is associated with a $f$-value estimating how much utility will be attained if we expand $s_{t}$:
@@
where $g(s_{t})$ denotes the aggregated utility from the initial state $s_{1}$; $h(s_{t})$ is the heuristic value for measuring the probability of reaching the correct answer derived from $s_{t}$; $\lambda$ is a coefficient to balance the importance of $g(s_{t})$ and $h(s_{t})$ terms.
@@
Since enumerating all possible next reasoning steps is intractable, in practice we restrict the alternatives to the top-K of all step candidates returned by LLM, and thus Eq. (6) can be written as $f(s_{t})=g(s_{t})+\lambda\max_{a_{t}\in\text{top-K}(\pi_{\theta}(\cdot|s_{t}))}Q^{*}(s_{t},a_{t})$.
@@ 4.1 Estimation of Optimal Q-value
where $s_{t}=[q_{i};a_{1};\dots;a_{t-1}]$ is the partial reasoning trace by timestep $t-1$ in $\mathbf{a}_{i_{j}}$ and $\hat{y}(s_{t},a_{t})$ is the label that approximates the true optimal Q-value, specifically $Q^{*}(s_{t},a_{t})$.
@@
Staring from the state-action pair $(s_{t},a_{t})$, we first perform random rollout or MCTS [27, 26] with policy $\pi_{\theta}$ for enough iterations, forming a trajectory pool $P$. After that, we use the best reasoning sequence with the highest accumulated rewards to construct the Q-value label:
@@ 4.2 Deliberative Planning with A*
Once obtaining the proxy Q-value model $\hat{Q}$, we can plug it to Eq. (6) to compute the $f$-value of each state and perform best-first search with A*. Alg. 1 illustrates the deliberative planning process.
@@
2:$unvisited\leftarrow\{q\}$, $visited\leftarrow\emptyset$
3:while $unvisited\neq\emptyset$ do
4:     $s\leftarrow\arg\max_{s^{\prime}\in unvisited}f(s^{\prime})$
5:     $unvisited\leftarrow unvisited\backslash\{s\}$, $visited\leftarrow visited\cup\{s\}$
6:     if $s$ is a terminal state then return the complete trajectory $s$
@@
9:         if $s^{\prime}\notin visited$ then $unvisited\leftarrow unvisited\cup\{s^{\prime}\}$
Specifically, we maintain a set for storing state candidates to be explored, denoted as $unvisited$, which initially only contains the input question $q$, and another set $visited$ to record the visited states. Each step we pick the state $s$ with the maximum $f$-value from the set $unvisited$ and expand it by querying the top-K alternatives with the LLM policy $\pi_{\theta}$. After that, both $visited$ and $unvisited$ sets will be updated and this process repeats until the terminal state (a complete trajectory) is reached. Finally, we extract the answer part of the terminal state $s$ as the result.